# A new hierarchy for complex plane curves

### v2

@@ Abstract.
We define the type of a plane curve as the initial degree of the corresponding Bourbaki ideal. Then we show that this invariant behaves well with respect to the union of curves. The curves of type 00 are precisely the free curves and the curves of type $1$ are the plus-one generated curves. In this note we show first that line arrangements and conic-line arrangements can have all the theoretically possible types. In the second part we study the properties of the curves of type $2$ and construct families of line arrangements and conic-line arrangements of this type.
@@ 1. Introduction
Let $S=\mathbb{C}[x,y,z]$ be the polynomial ring in three variables $x,y,z$ with complex coefficients, and let $C:f=0$ be a reduced curve of degree $d\geq 3$ in the complex projective plane $\mathbb{P}^{2}$. We denote by $J_{f}$ the Jacobian ideal of $f$, i.e. the homogeneous ideal in $S$ spanned by the partial derivatives $f_{x},f_{y},f_{z}$ of $f$, and by $M(f)=S/J_{f}$ the corresponding graded quotient ring, called the Jacobian (or Milnor) algebra of $f$. Consider the graded $S$-module of Jacobian syzygies of $f$, namely
@@
We say that $C:f=0$ is an $m$-syzygy curve if the module $F_{2}$ has rank $m$. Then the module $D_{0}(f)$ is generated by $m$ homogeneous syzygies, say $r_{1},r_{2},...,r_{m}$, of degrees $d_{j}=\deg r_{j}$ ordered such that
$1\leq d_{1}\leq d_{2}\leq...\leq d_{m}.$
We call these degrees as the exponents of the curve $C$, and $r_{1},...,r_{m}$ as a minimal set of generators for the module $D_{0}(f)$. Note that $d_{1}=mdr(f)$ is the minimal degree of a non-trivial Jacobian relation in $D_{0}(f)$. It is known that
@@
where ${\rm Sing}(C)$ denotes the set of all singular points of $C$ and $\tau_{p}$ is the Tjurina number of the singularity $(C,p)$.
@@
with $c_{1}\leq...\leq c_{m-2}$ and $d_{1}\leq...\leq d_{m}$. It follows from [20, Lemma 1.1] that one has
@@
for $j=1,...,m-2$ and some integers $\varepsilon_{j}\geq 1$. Using [20, Formula (13)], it follows that one has
(1.7) $d_{1}+d_{2}=d-1+\sum_{i=1}^{m-2}\varepsilon_{j}.$
This formula implies the following result, for details see [17, Theorem 2.3].
@@
Let $C:f=0$ be a reduced plane curve of degree $d$ and let $d_{1}$ and $d_{2}$ be the minimal degrees of a minimal system of generators for the module of Jacobian syzygies $D_{0}(f)$ as above. Then the following hold.
@@
Let $C:f=0$ be a reduced plane curve of degree $d$ and let $d_{1}$ and $d_{2}$ be the two smallest degrees of a minimal system of generators for the module of Jacobian syzygies $D_{0}(f)$ as above. Then we say that $C$ has type $t(C)$, where the positive integer $t(C)$ is defined by the formula
@@
The fact that this invariant plays a key role in the study of plane curves follows from the following alternative definition of it. Let $C:f=0$ be a reduced curve in $\mathbb{P}^{2}$ and $r_{1}$ be a minimal degree generator of the graded $S$-module $D_{0}(f)$. Then the corresponding Bourbaki ideal $B(C,r_{1})\subset S$ of $D_{0}(f)$ was introduced in [18, Section 5], and [17, Section 3], after being considered in a non-explicit way by du Plessis–Wall in [19] and the first author in [6]. Its properties were further explored in the recent paper [22]. It follows from the definition of the Bourbaki ideal $B(C,r_{1})$ that the type of the curve $C$ is just the initial degree of the homogeneous ideal $B(C,r_{1})$, namely
@@
Note that a curve $C$ is free exactly when $t(C)=0$, and a curve $C$ is plus-one generated exactly when $t(C)=1$. The type of a curve is a positive integer bounded by $22d-$ 2 in view of (1.2), and this upper bound is attained for instance by a smooth curve. For line arrangements the corresponding upper bound is $42d-$ 4 in view of (1.3), see Corollaries 1.9 and 1.11 below for more on this subject.
@@
Let $C_{1}:f_{1}=0$ and $C_{2}:f_{2}=0$ be two curves in $\mathbb{P}^{2}$, with no irreducible component in common. Then the union curve $C=C_{1}\cup C_{2}:f=f_{1}f_{2}=0$ satisfies
@@
If $C_{1}$ is free and $C_{2}$ is a line, then this result is already known, see [8, 23].
@@
Remark 1.6.
Consider the Fermat curve $C_{1}:f_{1}=x^{d}+y^{d}+z^{d}=0$ and the line arrangement $C_{2}:f_{2}=x^{d}+y^{d}=0$. Then the curve $C_{1}$ has type $d-1$ since the exponents of $C_{1}$ are $d_{1}=d_{2}=d_{3}=d-1$. On the other hand, the union $C=C_{1}\cup C_{2}$ has type 0, since $C$ is a free curve, see [11, Corollary 1.6]. It follows that there is no interesting lower bound for the type of $C=C_{1}\cup C_{2}$, without involving more information on $C_{1}$ and $C_{2}$.
@@
Let $C_{1}:f_{1}=0$ be a curve and let $d_{1}\leq d_{2}$ be the two smallest degrees of a minimal system of generators for the module of Jacobian syzygies $D_{0}(f_{1})$. Let $C_{2}:f_{2}=0$ be a smooth curve, such that the intersection $C_{1}\cap C_{2}$ consists of $\deg C_{1}\deg C_{2}$ points. If
@@
then the union curve $C=C_{1}\cup C_{2}:f=f_{1}f_{2}=0$ satisfies
@@
$30\leq t({\mathcal{A}})\leq\deg({\mathcal{A}})-$ 3 .
Conversely, for any integers $t$ and $d$ such that $30\leq t\leq d-$ 3, there are line arrangements in $\mathbb{P}^{2}$ such that $t({\mathcal{A}})=t$ and $\deg({\mathcal{A}})=d$.
To prove a result similar to Corollary 1.9 for conic-line arrangements is more complicated. The smooth conic $C_{1}$ in Example 1.8 satisfies $t(C_{1})=d-1$, while the conic-line arrangement $C$ from the same example satisfies $t(C)=d-2$. It is easy to check that 2 smooth conics meeting in 4 points give rise to conic-line arrangement ${\mathcal{C}}$ with $t({\mathcal{C}})=2=d-2$, since $d=\deg({\mathcal{C}})=4$. We believe that such examples are exceptional, and ask the following question.
Question 1.10.
@@
$30\leq t({\mathcal{C}})\leq\deg({\mathcal{C}})-$ 3 .
Conversely, for any integers $t$ and $d$ such that $30\leq t\leq d-$ 3, there are conic-line arrangements ${\mathcal{C}}$ containing at least one conic in $\mathbb{P}^{2}$ such that $t({\mathcal{C}})=t$ and $\deg({\mathcal{C}})=d$.
@@
The type of a line arrangement is not determined by the combinatorics. Indeed, the pair of arrangements
@@
$(x+5y+5z)(7x-4y-z)=0,$
constructed by Ziegler [32] and Yuzvinski [31], have the same combinatorics and as exponents $(5,6,6,6)$ and respectively $(6,6,6,6,6,6)$. It follows that the type of ${\mathcal{A}}$ is 3, and the type of ${\mathcal{A}}^{\prime}$ is 4.
One can make a loose analogy of the type of curves introduced in Definition 1.2 with Arnold’s notion of modality of function germs. In the case of function germs, the 00-modal (alias the simple singularities), the $1$-modal (alias the unimodal singularities) and the $2$-modal singularities are classified, see for instance [2], and the complexity of the corresponding geometry, e.g. reflected in their Milnor lattices, is increasing as the modality increases.
@@
Let us look at the first examples of curves $C$ having type $2$.
Example 1.16.
$f_{1}=(x^{2}+y^{2}-z^{2})(y-z)(x^{2}-z^{2})=0.$
$f_{1}=(-3x^{2}+xy+yz+xz)(-3y^{2}+xy+yz+zx)(-3z^{2}+xy+yz+zx)=0.$
This arrangement is known to be free with exponents $(2,3)$ and it has $3$ singularities of type $A_{5}$ and just one ordinary triple point $D_{4}$. If we add the conic $Q:f_{2}=x^{2}+y^{2}+z^{2}=0$, we get new conic arrangement $\mathcal{C}$ which has type $2A$ and exponents $(4,5,6)$. Note that the conic $Q$ is generic with respect to the arrangement $\mathcal{C}_{1}$, namely the intersection $\mathcal{C}_{1}\cap Q$ consists of $12$ points, which give rise to $12$ nodes for the arrangement $\mathcal{C}$
Remark 1.17.
One very natural way to obtain line arrangements of type $2$ is the following. Start with a free line arrangement ${\mathcal{A}}$ and let ${\mathcal{A}}_{i,j}$ be the line arrangement obtained from ${\mathcal{A}}$ by deleting two lines, namely $L_{i}$ and $L_{j}$. Then [4, Theorem 1.8] gives sufficient condition such that $t({\mathcal{A}}_{i,j})=2$. More precisely, with the notation from this result, we have the following possibilities. The arrangements in [4, Theorem 1.8] (1) have type $2B$, and the same holds for the arrangements in case (2), when $c_{1}<c_{2}$. On the other hands, the arrangements in [4, Theorem 1.8] (2) have type $2A$ if $c_{1}=c_{2}$ and $t({\mathcal{A}}_{i,j})\neq 1$. Moreover, in [4, Example 4.5], the author produces a free line arrangement ${\mathcal{A}}$ of $11$ lines, such that the arrangement ${\mathcal{A}}_{3,4}$ has type $2A$ and the arrangements ${\mathcal{A}}_{2,11}$ and ${\mathcal{A}}_{3,5}$ have type $2B$.
@@
Example 1.20.
An illustration of Theorem 1.19 occurs in Example 1.16 (iii), where $(m_{1},m_{2})=(2,3)$. Note that in Example 1.16 (ii) the condition that the intersection $Q\cap{\mathcal{A}}^{\prime}$ consists of $102\deg{\mathcal{A}}^{\prime}=2(m_{1}+m_{2}+1)=$ 10 points is not fulfilled, since $\tau(\mathcal{CL})=24$ by Proposition 1.13, and hence
$\tau(\mathcal{CL}_{1})+10=22\neq\tau(\mathcal{CL}).$
This explains why the type $2B$ may occur here.
@@
Let ${\mathcal{B}}_{1}$ (resp. ${\mathcal{B}}_{2}$) be the arrangement consisting of $n_{1}$ lines passing through a point $a_{1}\in\mathbb{P}^{2}$ (resp. $n_{2}$ lines passing through a point $a_{2}\in\mathbb{P}^{2}\setminus{\mathcal{B}}_{1}$) and such that $1\leq n_{1}\leq n_{2}$ and $n_{1}+n_{2}>2$. Then the arrangement ${\mathcal{A}}^{\prime}={\mathcal{B}}_{1}\cup{\mathcal{B}}_{2}$ has type 00 and exponents $(1,n_{1}+n_{2}-2)$ for $n_{1}=1$, and type $1$ and exponents $(n_{1},n_{2},n_{1}+n_{2}-2)$ for $n_{1}\geq 2$.
@@
Let ${\mathcal{A}}^{\prime}$ be the line arrangement from Proposition 1.22 and assume that $3\leq n_{1}\leq n_{2}$. Let $L$ be a line such that the intersection $L\cap{\mathcal{A}}^{\prime}$ consists of $n_{1}+n_{2}-1$ points. Then the arrangement ${\mathcal{A}}={\mathcal{A}}^{\prime}\cup L$ has has type $2A$ with exponents of $(n_{1}+1,n_{2}+1,n_{1}+n_{2}-2)$.
Finally we present an enumerative description of line arrangements of type $2$. Let us first fix the notation. Let $\mathcal{L}\subset\mathbb{P}^{2}$ be an arrangement of $d$ lines. We denote by $t_{k}:=t_{k}(\mathcal{L})$ the number of $k$-fold intersection points of $\mathcal{L}$ and we define the maximal multiplicity of $\mathcal{L}$ as follows:
@@
Example 1.26.
$Q(x,y,z)=xyz(x+y-2z)(x-3y+z)(-5x+y+z)(x+y+z).$
@@ 2. Some preliminaries
Let $C\,:f=0$ be a reduced plane curve of degree $d$ and $d_{1}={\rm mdr}(C)$. Then the following hold.
@@
If we set $T=3(d-2)$, then the sequence $n(f)_{k}$ is symmetric with respect to the middle point $T/2$, that is one has
@@
(2.2) $0\leq n(f)_{0}\leq n(f)_{1}\leq\ldots\leq n(f)_{[T/2]}\geq n(f)_{[T/2]+1}\geq\ldots\geq n(f)_{T}\geq 0$
Moreover, for a $3$-syzygy curve $C$ of degree $d$ with exponents $(d_{1},d_{2},d_{3})$, we have the following formula for the initial degree of the graded module $N(f)$, see [17, Theorem 3.9],

### v3

@@ Abstract.
We define the type of a plane curve as the initial degree of the corresponding Bourbaki ideal. Then we show that this invariant behaves well with respect to the union of curves. Curves of type 0 are precisely the free curves, while curves of type $1$ are the plus-one generated curves. In this paper, we first show that line arrangements and conic-line arrangements can exhibit all the theoretically possible types. In the second part, we study the properties of the curves of type $2$ and construct families of line arrangements and conic-line arrangements of this type.
@@ 1. Introduction
Let $S=\mathbb{C}[x,y,z]$ be the polynomial ring in three variables $x,y,z$ with complex coefficients, and let $C:f=0$ be a reduced curve of degree $d\geq 3$ in the complex projective plane $\mathbb{P}^{2}$. The ring $S$ and all the graded $S$-modules considered below are graded with respect to the natural grading. We denote by $J_{f}$ the Jacobian ideal of $f$, i.e. the homogeneous ideal in $S$ spanned by the partial derivatives $f_{x},f_{y},f_{z}$ of $f$, and by $M(f)=S/J_{f}$ we denote the corresponding graded quotient ring, called the Jacobian (or Milnor) algebra of $f$. Consider the graded $S$-module of Jacobian syzygies of $f$, namely
@@
We say that $C:f=0$ is an $m$-syzygy curve if the module $F_{2}$ has rank $m$. Then the module $D_{0}(f)$ is generated by $m$ homogeneous syzygies, say $r_{1},r_{2},\ldots,r_{m}$, of degrees $d_{j}=\deg r_{j}$ ordered such that
$1\leq d_{1}\leq d_{2}\leq\ldots\leq d_{m}.$
We call these degrees as the exponents of the curve $C$, and $r_{1},\ldots,r_{m}$ as a minimal set of generators for the module $D_{0}(f)$. Note that $d_{1}=mdr(f)$ is the minimal degree of a non-trivial Jacobian relation in $D_{0}(f)$. It is known that
@@
where ${\rm Sing}(C)$ denotes the set of all singular points of $C$, and $\tau_{p}$ is the Tjurina number of the singularity $(C,p)$.
@@
with $c_{1}\leq\ldots\leq c_{m-2}$ and $d_{1}\leq\ldots\leq d_{m}$. It follows from [20, Lemma 1.1] that one has
@@
for $j=1,\ldots,m-2$ and some integers $\varepsilon_{j}\geq 1$. Using [20, Formula (13)], it follows that one has
(1.7) $d_{1}+d_{2}=d-1+\sum_{j=1}^{m-2}\varepsilon_{j}.$
This formula implies the following result, see [17, Theorem 2.3] for details.
@@
Let $C:f=0$ be a reduced plane curve of degree $d$ and let $d_{1}$ and $d_{2}$ be the minimal degrees of a minimal system of generators for the module of Jacobian syzygies $D_{0}(f)$ as above. Then the following conditions hold:
@@
Let $C:f=0$ be a reduced plane curve of degree $d$ and let $d_{1}$ and $d_{2}$ be the two smallest degrees of a minimal system of generators for the module of Jacobian syzygies $D_{0}(f)$ as above. We then say that $C$ has type $t(C)$, where $t(C)$ is a positive integer defined by the formula
@@
The fact that this invariant plays a key role in the study of plane curves follows from the following alternative definition. Let $C:f=0$ be a reduced curve in $\mathbb{P}^{2}$ and $r_{1}$ be a minimal degree generator of the graded $S$-module $D_{0}(f)$. Then the corresponding Bourbaki ideal $B(C,r_{1})\subset S$ of $D_{0}(f)$ was introduced in [18, Section 5], and [17, Section 3], after being considered in a non-explicit way by du Plessis and Wall in [19], and the first author in [6]. Its properties were further explored in the recent paper [22]. It follows from the definition of the Bourbaki ideal $B(C,r_{1})$ that the type of the curve $C$ is just the initial degree of the homogeneous ideal $B(C,r_{1})$, namely
@@
Recall that the Bourbaki degree $\operatorname{Bour}(C)$ of the curve $C$ is defined by
(1.9) $\operatorname{Bour}(C)=\deg B(C,r_{1})$
and it satisfies the equality
(1.10) $\operatorname{Bour}(C)=(d-1)^{2}-d_{1}(d-d_{1}-1)-\tau(C),$
see [18, Theorem 5.1 (1)] and [22, Theorem 2.1]. It follows that, if $d_{1}\leq(d-1)/2$, then
(1.11) $\operatorname{Bour}(C)=\nu(C),$
see Theorem 2.1 below.
Note that a curve $C$ is free exactly when $t(C)=0$ and also if and only if $\operatorname{Bour}(C)=0$. On the other hand, a curve $C$ is plus-one generated exactly when $t(C)=1$. A plus-one generated curve is a 3-syzygy curve with exponents $(d_{1},d_{2},d_{3})$ satisfying $d_{1}+d_{2}=d$, see [17, Theorem 2.3 (2)]. For such a curve one has
(1.12) $\tau(C)=(d-1)^{2}-d_{1}(d-d_{1}-1)+d_{2}-d_{3}-1,$
see [17, Proposition 2.1 (4)]. It follows that, for a plus-one generated curve $C$ with exponents $(d_{1},d_{2},d_{3})$ one has
(1.13) $\operatorname{Bour}(C)=d_{3}-d_{2}+1,$
and hence $\operatorname{Bour}(C)=1$ if and only if $d_{2}=d_{3}$. This occurs when $C$ is a nearly free curve, which is a special type of plus-one generated curve.
The type of a curve is a positive integer bounded by 2d22d-2 in view of (1.2), and this upper bound is attained, for instance, by a smooth curve. For line arrangements the corresponding upper bound is 2d42d-4 in view of (1.3), see Corollaries 1.8 and 1.10 below for more on this subject.
@@
Let $C_{1}:f_{1}=0$ and $C_{2}:f_{2}=0$ be two curves in $\mathbb{P}^{2}$ with no irreducible component in common. Then the union curve $C=C_{1}\cup C_{2}:f=f_{1}\cdot f_{2}=0$ satisfies
@@
Remark 1.5.
If $C_{1}$ is free and $C_{2}$ is a line, then this result is already known – see [8, 23].
@@
Let $C_{1}:f_{1}=0$ be a curve and let $d_{1}\leq d_{2}$ be the two smallest degrees of a minimal system of generators for the module of Jacobian syzygies $D_{0}(f_{1})$. Let $C_{2}:f_{2}=0$ be a smooth curve such that the intersection $C_{1}\cap C_{2}$ consists of $\deg C_{1}\cdot\deg C_{2}$ points. If
@@
then the union curve $C=C_{1}\cup C_{2}:f=f_{1}\cdot f_{2}=0$ satisfies
@@
$0\leq t({\mathcal{A}})\leq\deg({\mathcal{A}})-3.$
Conversely, for any integers $t$ and $d$ such that $0\leq t\leq d-3$, there are line arrangements in $\mathbb{P}^{2}$ such that $t({\mathcal{A}})=t$ and $\deg({\mathcal{A}})=d$.
Proving a result similar to Corollary 1.8 for conic-line arrangements is more complicated. The smooth conic $C_{1}$ in Example 3.4 satisfies $t(C_{1})=d-1$, while the conic-line arrangement $C$ from the same example satisfies t(C)=d2t(C)=d-2. It is easy to check that 2 smooth conics meeting in 4 points give rise to conic-line arrangement $\mathcal{C}$} with $t({\mathcal{C}})=2=d-2$, since $d=\deg({\mathcal{C}})=4$. We believe that such examples are exceptional, and ask the following question.
Question 1.9.
@@
$0\leq t({\mathcal{C}})\leq\deg({\mathcal{C}})-3.$
Conversely, for any integers $t$ and $d$ such that $0\leq t\leq d-3$, there are conic-line arrangements $\mathcal{C}$} containing at least one conic in $\mathbb{P}^{2}$ such that $t({\mathcal{C}})=t$ and $\deg({\mathcal{C}})=d$.
@@
Before presenting further results, we would like to make the following observation regarding the type.
@@
The type in the class of line arrangements is not determined by the combinatorics. Indeed, the pair of arrangements
@@
(x+5y+5z)(7x4yz)=0(x+5y+5z)(7x-4y-z)=0
constructed by Ziegler [32] and Yuzvinski [31] have the same combinatorics, but the exponents are $(5,6,6,6)$ and $(6,6,6,6,6,6)$, respectively. Therefore, the type of $\mathcal{A}$} is 33, while the type of $\mathcal{A}$$^{\prime}$ is 44.
One can make a loose analogy of the type of curves introduced in Definition 1.2 with Arnold’s notion of modality of function germs. In the case of function germs, the 0-modal (alias the simple singularities), the $1$-modal (alias the unimodal singularities) and the $2$-modal singularities are classified, see for instance [2], and the complexity of the corresponding geometry, e.g. reflected in their Milnor lattices, is increasing as the modality increases.
@@
Let ${\mathcal{B}}_{1}$ (resp. ${\mathcal{B}}_{2}$) be the arrangement consisting of $n_{1}$ lines passing through a point $a_{1}\in\mathbb{P}^{2}$ (resp. $n_{2}$ lines passing through a point $a_{2}\in\mathbb{P}^{2}\setminus{\mathcal{B}}_{1}$) and such that $1\leq n_{1}\leq n_{2}$ and $n_{1}+n_{2}>2$. Then the arrangement ${\mathcal{A}}^{\prime}={\mathcal{B}}_{1}\cup{\mathcal{B}}_{2}$ has type 0 and exponents $(1,n_{1}+n_{2}-2)$ for $n_{1}=1$, and type $1$ and exponents $(n_{1},n_{2},n_{1}+n_{2}-2)$ for $n_{1}\geq 2$.
@@
Let $\mathcal{A}$$^{\prime}$ be the line arrangement from Proposition 1.18 and assume that $3\leq n_{1}\leq n_{2}$. Let $L$ be a line such that the intersection $L\cap{\mathcal{A}}^{\prime}$ consists of $n_{1}+n_{2}-1$ points. Then the arrangement ${\mathcal{A}}={\mathcal{A}}^{\prime}\cup L$ has has type $2A$ with exponents $(n_{1}+1,n_{2}+1,n_{1}+n_{2}-2)$.
Finally, we present an enumerative description of line arrangements of type $2$. Let us first fix the notation. Let $\mathcal{L}\subset\mathbb{P}^{2}$ be an arrangement of $d$ lines. We denote by $t_{k}:=t_{k}(\mathcal{L})$ the number of $k$-fold intersection points of $\mathcal{L}$, and we define the maximal multiplicity of $\mathcal{L}$ as follows:
@@
The content of this paper is outlined below. Section 2 provides a recap of some necessary preliminaries. The subsequent sections then proceed to prove these results and provide a number of useful examples and related additional remarks.
@@ 2. Some preliminaries
Let C:f=0C\,:f=0 be a reduced plane curve of degree $d$ and $d_{1}={\rm mdr}(C)$. Then the following hold.
@@
If we set T=3(d2)T=3(d-2), then the sequence $n(f)_{k}$ is symmetric with respect to the middle point $T/2$, that is one has
@@
(2.2) $0\leq n(f)_{0}\leq n(f)_{1}\leq\ldots\leq n(f)_{[T/2]}\geq n(f)_{[T/2]+1}\geq\ldots\geq n(f)_{T}\geq 0.$
Moreover, for a $3$-syzygy curve $C$ of degree $d$ with exponents $(d_{1},d_{2},d_{3})$, we have the following formula for the initial degree of the graded module $N(f)$, see [17, Theorem 3.9]: